Determine the critical values for the confidence interval for the population standard deviation from the given values. Round your answers to three decimal places. n=21 and α=0.1 .
Answer #1

We have given,              
              
Sample size =n=21      
Degree of freedom =n-1=21-1=20  
Level of significance=0.1      
              
\chi_{L}^{2}= 10.851 by using chi square critical value table or Excel =CHIINV(1-(0.1/2),20)
\chi_{U}^{2}= 31.41 by using chi square critical value table or Excel =CHIINV((0.1/2),20)